### 11 Mar 2026 v1

@@ 2.5 PAPC
The PAPC (Parallel Architectures, Programming, and Compilers) Group at the University of Malaga (UMA) conducts research at the intersection of parallel architectures, compilers, and high-performance programming, with a special focus on optimization, performance, and scalability of scientific and machine learning software. In recent years, it has incorporated a growing line of research into quantum technologies and quantum machine learning, where it applies its expertise in performance engineering and parallel models to evaluate, accelerate, and make quantum-classical hybrid workflows more reproducible.
The research group presents the work entitled “Redes tensoriales aplicadas a redes neuronales cuánticas: evaluación de eficiencia y rendimiento[1], which presents a systematic evaluation of the use of tensor networks (TNs) (such as MPS/PEPS) to simulate and train quantum and hybrid neural networks in the NISQ era. The work analyzes how TNs allow quantum states and operations to be emulated in a compact way, reducing computational resources compared to dense simulation, and studies the trade-off between efficiency and accuracy as entanglement increases. The results show that, for specific architectures and tasks, TNs can accelerate development and experimentation in QML without always relying on quantum hardware, while pointing out limits and best practices (choice of TN architecture, bond dimension control, and variational circuit design) to maintain predictive quality at a reasonable computational cost.

### 16 Mar 2026 v2

@@ 2.5 PAPC
The PAPC (Parallel Architectures, Programming, and Compilers) Group at the University of Malaga (UMA) conducts research at the intersection of parallel architectures, compilers, and high-performance programming, with a special focus on the optimization, performance, and scalability of scientific and machine learning software. In recent years, the group has also developed a growing and ambitious research line in quantum technologies and quantum machine learning, applying its expertise in performance engineering and parallel computing to accelerate and improve the reproducibility of hybrid quantum-classical workflows, enable large-scale simulation of quantum computation using HPC resources, and develop new quantum algorithms for the NISQ era. This group is also actively involved in the Quant·UMA initiative.
The research group presents the work entitled “Redes tensoriales aplicadas a redes neuronales cuánticas: evaluación de eficiencia y rendimiento[1], which provides a systematic evaluation of tensor networks (TNs), as feature extractors placed before the variational quantum circuit (VQC) within a hybrid quantum neural network (HQNN) architecture. The study analyzes how TNs can improve the alignment of input data with the quantum block through a quantum-inspired processing scheme. Early results suggest that, for specific architectures and tasks, TNs can transform data using fewer parameters and prepare it more effectively for the quantum circuit, while also highlighting limitations and best practices—such as the choice of TN architecture, bond-dimension control, and variational circuit design—to maintain predictive performance at a reasonable computational cost.